OpenStudy (anonymous):

The density of a mixture of florin and chlorine gases is 1.77g/L at 14 degree celcius at 0.893 atm. Calculate the mass percent of the gases.

OpenStudy (jfraser):

find the density of each pure gas using the ideal gas law, then set up a weighted average. Since\[P*V = n*R*T\]then\[d = \frac{g}{V} = \frac{MM*P}{R*T}\]